Grade 9Mechanics

A cuboid has dimensions of 0.4 m ×0.6 m×0.2m and the weight of 288 kg-d. What is the maximum pressure exerted by the cuboid?(Ans-35280 Pa)

chetan
8 Years agoGrade 9

1 Answer

Arun
8 Years ago
Dear Chetan
 
Maximum pressure = Force/ min Area
 = 288 * 9.8 /(0.4 * 0.2) = 2822.4 /0.08
 
 = 35280 N /m² = 35280 Pascals